Vaughn Company reported the following amounts in the stockholders equity section of its December 31, 2019, balance sheet. PrNo. Account Titles and Explanation Debit Credit 1. Dividends Payable - Preferred 22000 Dividends Payable - Common 37000 CashTreasury Stock 24000 Paid-in Capital from Treasury Stock 8100 4. Cash 55650 Preferred Stock 53000 Paid-in Capital in Excess oVAUGHN COMPANY Stockholders Equity December 31, 2020 Capital Stock Common Stock 101100 Preferred Stock 253000 Total Capital
Answer #1
Journal General
No. Account Titles and Explanation Debit Credit
      1 Dividends Payable - Preferred     22,000
Dividends Payable - Common     37,000
Cash 59,000
[dividend paid]
      2 Treasury stock     76,000
Cash 76,000
[shares reupurchased]
      3 Land     32,100
Treasury stock 24,000
Paid-in Capital from Treasury Stock     8,100
[treasury stock reissued]
      4 Cash     55,650
Preferred stock 53,000
Paid-in Capital in Excess of Par - Preferred Stock     2,650
      5 Retained earnings     80,840
Common Stock Dividend Distributable     8,600
Paid-in Capital in Excess of Par - Common Stock 72,240
[dividend declared]  
      6 Common Stock Dividend Distributable       8,600
Common Stock     8,600
      7 Retained earning     65,670
Dividends Payable - Preferred 27,830
Dividends Payable - Common 37,840
[dividend declared]
Vaughn Company
Stockholders' Equity
December 31. 2020.
Capital Stock
Common Stock 101,100
Preferred Stock 253,000
Total Capital Stock 354,100
Additional Paid-in Capital     82,990
Total Paid-in Capital 437,090
Retained Earnings 266,490
Total Paid-in Capital and Retained Earnings 703,580
Less: Treasury Stock     52,000
Total Stockholders' Equity 651,580

Please find below table useful to compute desired results: -

B E F G Particulars Given =11/100 100 Preferred Stock Common Stock Additional Paid-in capital Retained Earnings 10000 2000 =F

C =B11*c11 294 30 31 B Cash Preferred stock Paid-in Capital in Excess of Par - Preferred Stock =B11*D2 =C29-D30 =(F3-B9+B10)*

End results would be as follows: -

D E F G Particulars Given 11% 100 5 10,000 92,500 2.000 18,500 Preferred Stock Common Stock Additional Paid-in capital Retain

D E F G A 4 C 55,650 B Cash Preferred stock Paid-in Capital in Excess of Par - Preferred Stock 53,000 2.650 80.840 8,600 72.2